300 mm 300 mm Determine the internal forces at Point J when a=0. lec 480 mm 240 mm С D 780 N
Answer #1

so 300 mm А EMA=0 (5) -A By BO 780X0.720 t By (0-3) =( 720 mA B J S 1872 N (+) By 730N { Fy so (1+) А (13) - 1872 = 0 A = 202